OpenStudy (anonymous):

I am suppose to find the slope, y-intercept, and graph it, x+4=5

OpenStudy (anonymous):

would I make it 4=-x+5?

terenzreignz (terenzreignz):

Are you sure that there's no y?

OpenStudy (anonymous):

positive

terenzreignz (terenzreignz):

Well then, here's a nice shortcut. If there's no y, then the slope is AUTOMATICALLY infinite or it does not exist. Because it'd be a vertical line. And you know that the slope of a vertical line is undefined.

OpenStudy (anonymous):

Are you sure from the equation ? \[x+4=5 ??\]

OpenStudy (anonymous):

yes @Noura11

OpenStudy (anonymous):

@terenzreignz then what's the y-intercept?

OpenStudy (anonymous):

and how do I graph it?

terenzreignz (terenzreignz):

Well, @Noura11 maybe you were intending to give a more detailed explanation? :) By all means...

OpenStudy (anonymous):

First , when we have an equation of the line like this : \[x=a\] then , there is no slope, or we can say that the slope is infinitive. This line will be a vertical line ! To draw a vertical line we put the point (a,0) and trace a vertical line passing by that point !

OpenStudy (anonymous):

what about the y-intercept

OpenStudy (anonymous):

it's 0?

OpenStudy (anonymous):

and where along x do I draw a vertical line?

OpenStudy (anonymous):

No ! If the equation is : \[x=a\] and : \[a\ne0\] the there is no y-intercept. But if the equation is : \[x=0\] then 0 is the y-intercept

OpenStudy (anonymous):

so I put no y-intercept,slope=undefined, but where along x do I put the vertical line

terenzreignz (terenzreignz):

Simple... solve the equation for x, and that's where you put the vertical line.

OpenStudy (anonymous):

@terenzreignz is that right?

terenzreignz (terenzreignz):

So far, yes. You just have to figure out where to put your line, and that's where solving for x comes in.

OpenStudy (anonymous):

x=1

terenzreignz (terenzreignz):

Correct :)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

@terenzreignz what if the equation was y=-2?

OpenStudy (anonymous):

y=mx+b, so would m and x be 0 while b=-2?

terenzreignz (terenzreignz):

Yeah, actually, yes. You've answered your own question LOL Remember how I said that if there's no mention of y, then the slope is undefined? On the flipside, if there's no mention of x, then the slope is automatically zero. :)

OpenStudy (anonymous):

so slope=0 and y-intercept=-2

OpenStudy (anonymous):

how do I graph a slope of 0? @terenzreignz

terenzreignz (terenzreignz):

Undefined slope = vertical Zero slope = horizontal Graph the horizontal line passing through b (your y-intercept)

OpenStudy (anonymous):

oh got it thanks!